OpenStudy (anonymous):

A system of equations is shown below: x + y = 3 2x – y = 6 The x-coordinate of the solution to this system of equations is _____. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (anonymous):

@ganeshie8 its only 1 question... help please?

OpenStudy (anonymous):

first subtract the y's from the equation

OpenStudy (anonymous):

sub 1st equation from the 2nd -x=-3 x=3

OpenStudy (anonymous):

apply x value in the first equation 3+y=3 y=3-3 y=0

OpenStudy (anonymous):

now you can check by applying x and y value in the 2nd equation 2(3)-0=6 6=6

OpenStudy (anonymous):

So... the x coordination is (-3,3) right??

OpenStudy (anonymous):

no x=3

OpenStudy (anonymous):

from where this -3 came

OpenStudy (anonymous):